# Sequences with identical autocorrelation functions

### v2

@@ 4. Equicorrelationality of generalized palindromes
Let $F$ be a subfield of ${\mathbb{C}}$ and let $f$ and $g$ be generalized palindromes with $\overline{f(z)}=uz^{j}f(z)$ and $\overline{g(z)}=vz^{k}g(z)$ for some $u,v\in{\mathbb{C}}^{\times}$ and $j,k\in{\mathbb{Z}}$. Suppose that $f,g\in F[z,z^{-1}]$ and that $f$ and $g$ are equicorrelational. Then either (i) $f=g=0$ or else (ii) $v/u=\overline{w}/w$ for some $w\in F^{\times}$ and $j\equiv k\pmod{2}$.
@@ 5. Unequivocal integers
We let $s=\gcd(a,c)$ and $t=\gcd(b,d)$; these are nonzero because $a,b,c,d$ are nonzero. Then we define $\alpha,\beta,\gamma,\delta$ as in (ii), so that $a=s\alpha$, $b=t\beta$, $c=s\gamma$, and $d=t\delta$. We note that $[f]=[g]$ if and only if there is a $u\in{\mathbb{C}}[z,z^{-1}]^{\times}$ such that $f=ug$. So $[f]=[g]$ if and only if there is a $u\in{\mathbb{C}}[z,z^{-1}]^{\times}$ such that $u(z)s(z^{m})\alpha(z^{m})t(z^{n})\beta(z^{n})=s(z^{m})\gamma(z^{m})t(z^{n})\delta(z^{n})$. Notice that $s(z^{m})$ and $t(z^{n})$ are nonzero because $s(z)$, $t(z)$, $m$, and $n$ are all nonzero. Since ${\mathbb{C}}[z,z^{-1}]^{\times}$ is an integral domain, this means that $[f]=[g]$ if and only if there is a $u\in{\mathbb{C}}[z,z^{-1}]^{\times}$ such that $u(z)\alpha(z^{m})\beta(z^{n})=\gamma(z^{m})\delta(z^{n})$. But $\alpha(z^{m})$ is relatively prime to $\gamma(z^{m})$ and $\beta(z^{n})$ is relatively prime to $\delta(z^{n})$ by Lemma 2.4, and ${\mathbb{C}}[z,z^{-1}]^{\times}$ is a unique factorization domain (since it is a Euclidean domain). Thus, $[f]=[g]$ if and only if both $[\alpha(z^{m})]=[\delta(z^{n})]$ and $[\beta(z^{n})]=[\gamma(z^{m})]$. ∎
@@
Lemma 5.4 applies to sequences $a$, $b$, $c$, and $d$ that need not be binary (nor, more generally, $k$-ary for some $k$), and even if these four sequences are binary (or $k$-ary), the parameters $m$ and $n$ might be such that the combined sequences $f$ and $g$ are not be binary (or $k$-ary). In many practical scenarios, we would want to constrain $a$, $b$, $c$, $d$, $m$, and $n$ so as to produce binary (or $k$-ary) $f$ and $g$, and we examine such situations in the following result, which makes use of 5.1.
@@ Acknowledgments
The authors thank Showmic Islam, Christina Koch, Andrew Owen, and Mats Rynge of the Open Science Grid for their assistance.

### v3

@@ 4. Equicorrelationality of generalized palindromes
Let $F$ be a subfield of ${\mathbb{C}}$ and let $f$ and $g$ be generalized palindromes with $\overline{f}=uz^{j}f$ and $\overline{g}=vz^{k}g$ for some $u,v\in{\mathbb{C}}^{\times}$ and $j,k\in{\mathbb{Z}}$. Suppose that $f,g\in F[z,z^{-1}]$ and that $f$ and $g$ are equicorrelational. Then either (i) $f=g=0$ or else (ii) $v/u=\overline{w}/w$ for some $w\in F^{\times}$ and $j\equiv k\pmod{2}$.
@@ 5. Unequivocal integers
We let $s=\gcd(a,c)$ and $t=\gcd(b,d)$; these are nonzero because $a,b,c,d$ are nonzero. Then we define $\alpha,\beta,\gamma,\delta$ as in (ii), so that $a=s\alpha$, $b=t\beta$, $c=s\gamma$, and $d=t\delta$. We note that $[f]=[g]$ if and only if there is a $u\in{\mathbb{C}}[z,z^{-1}]^{\times}$ such that $f=ug$. So $[f]=[g]$ if and only if there is a $u\in{\mathbb{C}}[z,z^{-1}]^{\times}$ such that $u(z)s(z^{m})\alpha(z^{m})t(z^{n})\beta(z^{n})=s(z^{m})\gamma(z^{m})t(z^{n})\delta(z^{n})$. Notice that $s(z^{m})$ and $t(z^{n})$ are nonzero because $s(z)$, $t(z)$, $m$, and $n$ are all nonzero. Since ${\mathbb{C}}[z,z^{-1}]^{\times}$ is an integral domain, this means that $[f]=[g]$ if and only if there is a $u\in{\mathbb{C}}[z,z^{-1}]^{\times}$ such that $u(z)\alpha(z^{m})\beta(z^{n})=\gamma(z^{m})\delta(z^{n})$. But $\alpha(z^{m})$ is relatively prime to $\gamma(z^{m})$ and $\beta(z^{n})$ is relatively prime to $\delta(z^{n})$ by Lemma 2.4, and ${\mathbb{C}}[z,z^{-1}]$ is a unique factorization domain (since it is a Euclidean domain). Thus, $[f]=[g]$ if and only if both $[\alpha(z^{m})]=[\delta(z^{n})]$ and $[\beta(z^{n})]=[\gamma(z^{m})]$. ∎
@@
Lemma 5.4 applies to sequences $a$, $b$, $c$, and $d$ that need not be binary (nor, more generally, $k$-ary for some $k$), and even if these four sequences are binary (or $k$-ary), the parameters $m$ and $n$ might be such that the combined sequences $f$ and $g$ are not binary (or $k$-ary). In many practical scenarios, we would want to constrain $a$, $b$, $c$, $d$, $m$, and $n$ so as to produce binary (or $k$-ary) $f$ and $g$, and we examine such situations in the following result, which makes use of 5.1.
@@ Acknowledgments
The authors thank Showmic Islam, Christina Koch, Andrew Owen, and Mats Rynge of the Open Science Grid for their assistance. The authors thank anonymous reviewers for encouragement and comments which helped them to improve the paper.